# Modified Schwarzschild spacetime

### v1

@@ Abstract
Considering an ansatz for $1+3+n$ dimensional static spacetime with uniform $n$-branes stacked in three-dimensional rotationally symmetric space, we found different classes of vacuum solutions of Einstein field equations. We analyzed these solutions through the Newtonian limit, Kretschmann scalar, and conserved energy from Landau–Lifshitz stress-energy pseudotensor. This revealed unusual physical properties for a class of solutions with nontrivial extension of the Schwarzschild spacetime with extra dimensions. Positive conserved energy is associated with repulsive gravity and physical singularity on the horizon, while negative conserved energy implies gravitational attraction and naked singularity. This sets the Schwarzschild solution in a broader context and provides new possible testing grounds for general relativity.
@@ 1 Introduction
Schwarzschild spacetime [2] is one of the most prominent examples of vacuum solutions of Einstein field equations. Since vacuum solutions correspond to zero Ricci tensor, they can be called Ricci-flat. The Schwarzschild solution can be interpreted as a black hole [3], and such objects are present in our Universe [4, 5, 6]. Black holes may generate Hawking radiation [7], providing a laboratory for testing quantum effects in gravity. Modification of the original Schwarzschild solution to an arbitrary number of space dimensions was first found by Tangherlini [8]. Rotating black holes with Kerr metric [9] can also be generalized to an arbitrary number of space dimensions. They are known as Myers–Perry black holes [10, 11]. There are also other different higher dimensional generalizations [12] including, for example, black rings [13].
@@
Physical motivation for extra dimensions originates mainly from Kaluza–Klein models [14, 15] attempting to solve hierarchy problem [16, 17] and string theory [18, 19, 20] developed as a promising candidate for a quantum theory of gravity. These theories usually require some kind of compactification of extra dimensions, which contradicts the aforementioned examples. Small extra dimensions can be present in a different class of objects known as black strings, p-branes or black branes [21, 22, 23, 24, 25, 26]. They may have black hole properties in macroscopic dimensions while containing compactified extra dimensions from String theory[27, 28, 29, 30]. Higher dimensional objects provide valuable testing grounds for studying quantum effects in general relativity[31, 32, 33].
@@
This paper provides an analysis of a special class of Ricci-flat spacetimes in $1+3+n$ dimensions under the assumption of spherical symmetry in the three-dimensional part and Euclidean symmetry in the $n$-dimensional part, $SO(3)\times E(n)$. Extra dimensions may be compactified into a flat $n$-torus with $S^{1}\times...\times S^{1}$ topology. The number of the standard space dimensions that cannot be compactified is fixed to three in this work since it is the physically most relevant case. Disregarding $1+3+n$ dimensional Minkowski spacetimes, there will be two different classes of solutions. The first class represents a simple extension of the $1+3$ dimensional Schwarzschild spacetime with the addition of $n$ flat dimensions. In contrast, the second class introduces a nontrivial form of $g_{AB}$ components of the spacetime metric, where indices $A$ and $B$ run through $n$ additional dimensions.
@@
The following section 2 provides vacuum solutions of Einstein field equations with two classes mentioned above, section 3 includes further analysis of properties of these spacetimes based on the Newtonian limit, Kretschmann scalar invariant and conserved energy from Landau–Lifshitz pseudotensor, and section 4 is dedicated to summary and discussion. The usual convention with the light speed set to unity, $c=1$, together with the spacetime metric signature (,+,+,)(-,+,+,...)( - , + , + , … ), will be used throughout this paper. The index 00 will be used for a time coordinate, lowercase Latin indices will indicate three space coordinates, $i=1,2,3$, and capital Latin indices will run through $n$ extra dimensions, $A=1,...,n$. All spacetime indices will be collectively denoted by Greek letters, $\mu=0,i,A$.
2 Ricci-flat solutions
@@
The existence of two different classes of Ricci-flat extensions of the Schwarzschild spacetime is obvious in the case with only one extra space dimension, often called black string. The metric of the trivial case can be written as
@@
where $t$ is time, $a$ is a constant, $r$ is radial coordinate, $d\Omega_{(2)}^{2}=d\vartheta^{2}+\sin^{2}\vartheta d\varphi^{2}$ is line element of a unit $2$-sphere, and $\zeta$ denotes coordinate in the additional dimension. Since the extra dimension is added most trivially, the spacetime with this metric is Ricci-flat like the original Schwarzschild spacetime. The Ricci tensor remains zero if this metric is rewritten through a formal change $\zeta\to it$, $t\to i\zeta$. The resulting metric reads
@@
Metrics (2.1) and (2.2) describe two spacetimes with different properties, which, for example, can be tested by a slowly moving particle far away from the center in the Newtonian limit. This simple example with only one extra space dimension can be generalized to the arbitrary number of extra dimensions, which will be examined in what follows.
Our goal is to find a metric of a Ricci-flat spacetime with three-dimensional spherical symmetry and $n$-dimensional Euclidean symmetry in extra dimensions, $SO(3)\times E(n)$, which can be interpreted as a stack of uniform $n$-branes in $1+3+n$ dimensions. We start with the ansatz
@@
where $\alpha$, $\beta$ and $\gamma$ are constants, and $f=f(r)$ is a function of only the radial coordinate. A more general ansatz may be parameterized by three independent functions of the radial coordinate [24], but in our analysis, we assume that these functions are given by different powers of one function $f(r)$. There are two reasons to start with the special choice of ansatz (2.3). A pragmatic reason is that with this choice, the condition of Ricci tensor being zero reduces to simple algebraic relations. The second reason is that one would expect only one potential event horizon, which occurs only when all components of the spacetime metric $g_{00}$, $g_{rr}$ and $g_{AB}$ are singular at the same radial coordinate, and the simplest way to achieve this is to consider our form of the ansatz. In principle, there may be Ricci-flat solutions with $SO(3)\times E(n)$ symmetry exhibiting more than one horizon; however, with the ansatz (2.3), we will not be able to find them. For example, authors of [34] consider a general form of ansatz with one additional dimension.
@@
$R_{00}=\alpha f^{\alpha-\beta}F_{1},\quad R_{rr}=F_{2},\quad R_{\vartheta\vartheta}=F_{3},$
$R_{\varphi\varphi}=F_{3}\sin^{2}\vartheta,\quad R_{AB}=-\gamma f^{\gamma-\beta}F_{1}\delta_{AB},$ (2.4)
@@
$F_{1}=\dfrac{1}{r}\dfrac{f^{\prime}}{f}+\dfrac{1}{4}\left(\alpha-\beta+n\gamma-2\right)\left(\dfrac{f^{\prime}}{f}\right)^{2}+\dfrac{1}{2}\dfrac{f^{\prime\prime}}{f},$
@@
$F_{3}=1-f^{-\beta}\left[1+\dfrac{1}{2}\left(\alpha-\beta+n\gamma\right)r\dfrac{f^{\prime}}{f}\right],$ (2.5)
with prime denoting differentiation with respect to the radial coordinate. All three functions $F_{1}$, $F_{2}$, and $F_{3}$ must be zero to have the Ricci-flat solution.
@@
where $a$ is an integration constant, and $q$ is given by constants from the ansatz. This solution is invalid if $\alpha-\beta+n\gamma=0$, and such case has to be treated separately. It is easy to see that in this case, the condition $F_{3}=0$ implies $f=1$, and we retrieve the simple $1+3+n$ dimensional Minkowski spacetime. The same is true if the integration constant $a$ is zero. By inserting (2.6) into expressions for $F_{1}$ and $F_{2}$ in (2.5) we find
@@
$\Phi=\dfrac{1}{\left(\alpha-\beta+n\gamma\right)^{2}}\dfrac{ar^{q-2}}{\left(1+ar^{q}\right)^{2}},$
@@
$C_{3}=2\beta^{2}-\alpha^{2}-n\gamma^{2}-\left(\alpha+n\gamma\right)^{2}.$ (2.8)
The problem of finding Ricci-flat solutions of the form (2.3) is now reduced to the system of algebraic equations.
First two conditions $C_{1}=0$ and $C_{2}=0$ are satisfied if $\alpha=-\beta-n\gamma$, which implies $q=-1$, and by inserting this expression for $\alpha$ into the third relation we also find
@@
The condition $C_{3}=0$ is then satisfied by two sets of possible solutions. One set corresponds to the choice $\gamma=0$, and for the second set, we have $02\beta+(n+1)\gamma=$ 0. With the choice $\gamma=0$ we have $\alpha=-\beta$. Components of the spacetime metric are then given by $f^{\alpha}=1+ar^{-1}$, $f^{\beta}=\left(1+ar^{-1}\right)^{-1}$, and $f^{\gamma}=1$, and the corresponding spacetime metric can be written as
@@
This is only a trivial extension of the original $1+3$ dimensional Schwarzschild solution, with metric (2.1) as a special case for $n=1$. A more interesting solutions correspond to the second choice with $02\beta+(n+1)\gamma=$ 0, where $\alpha/\beta=(n-1)/(n+1)$, and $\gamma/\beta=-2/(n+1)$. Spacetime metric then can be written as
@@
We can also check that for $n=1$ we retrieve (2.2). In the case with no extra dimensions, $n=0$, both sets of solutions (2.10) and (2.11) automatically reduce to the Schwarzschild solution. These solutions contain potential singularities. They may also contain horizons because vector field with components $\delta^{\mu}_{0}$ is a Killing vector in all of these spacetimes, and if $g_{00}=0$, the norm of such killing vector is zero. This will be discussed in more detail in the next section.
Different solutions similar to (2.11) can be found in the literature. Authors of [21] worked with low energy string theory action in ten dimensions, and they found a brane solution with a single Ramond-Ramond charge with nontrivial $g_{AB}$ components in terms of our notations. This has been studied further in many other works[35, 36, 37, 38, 39, 40, 41, 42]. Extra dimensions parameterized by coordinates $\zeta^{A}$ in (2.11) may be macroscopic with $\mathbb{R}^{n}$ topology, microscopic with $n$-torus topology, or a combination of both cases with some macroscopic and some microscopic dimensions. In all cases, there is Euclidean symmetry in extra dimensions, but one can assign extra dimensions to the part that is only spherically symmetric [43]. Note also that considering extra time dimensions instead of extra space dimensions, or having mixed time and space extra dimensions, would not change the results because changing signs of $g_{AB}$ components of the metric keeps Ricci tensor being zero. Compactification of the time dimension may also be considered, with implications for branes [44]. In the following section, where we will study some properties of the found solutions (2.10) and (2.11), we will assume the physically most relevant case with three macroscopic space dimensions with coordinates $x^{i}$ and $n$ compactified extra space dimensions with coordinates $\zeta^{A}$.
@@ 3 Physical properties
The Newtonian limit provides a very important insight into properties of branes containing spherically symmetric $1+3$ dimensional part [45, 46]. We will examine solutions (2.10) and (2.11) in this limit. To focus on physically most relevant cases, let us assume that all extra dimensions are microscopic with $S^{1}\times...\times S^{1}$ topology of a flat $n$-torus. Space is then $L$-periodic in $n$ space directions, $\zeta^{A}\in\left[0,L\right)$. Macroscopic objects with a size larger than $L$ then can move only in the $1+3$ dimensional section. The same is true for sufficiently long wavelength photons or particles with large enough de Broglie wavelengths. From the perspective of such objects, the spacetime in regions far from the origin, $r\gg|a|$, can be approximated as
@@
where $\phi$ is the Newtonian potential. Without any extra microscopic dimensions, it can be shown that $\psi=\phi$ in harmonic coordinates. The geodesic equation for a slowly moving object reduces to the Newtonian equation of motion, $\ddot{x}^{i}\approx-\phi_{,i}$, in this limit. In the case with trivial extension of the Schwarzschild spacetime (2.10), the form of the Newtonian potential is usual, $\phi=(a/2)r^{-1}$, and in the nontrivial case (2.11) we have
$\phi=-\dfrac{n-1}{n+1}\dfrac{a}{2}\dfrac{1}{r}=-\kappa M\dfrac{1}{r},$ (3.2)
where $M$ is the mass of a gravitating body in the context of classical mechanics, which is valid in the Newtonian limit. This provides the standard way of fixing the constant $a$ in the Schwarzschild solution as $a=-2\kappa M$. A consequence of a negative value of $a$ is the existence of a horizon. However, in the modified case with the spacetime metric (2.11), the sign is opposite, which implies naked singularities for positive masses $M$. Negative $a$ and the existence of horizon is then associated with negative mass $M$, which means repulsive gravity in the Newtonian limit. Other examples of different types of black holes with such property are studied, for example, in [47]. We may also notice that the presence of extra dimensions does not change the position of the horizon or its properties considerably. For example, a contrary case can be found in [48]. Note also that for $n=1$, the spacetime metric in the nontrivial case is (2.2), and the Newtonian potential vanishes, which means zero mass $M$.
In the case of the ordinary Schwarzschild metric, there is no physical singularity on the horizon. This can be easily revealed by the calculation of the Kretschmann scalar invariant defined through Riemann curvature tensor as $K=R^{\mu\nu\lambda\sigma}R_{\mu\nu\lambda\sigma}$. For the Schwarzschild spacetime it is $K_{\textrm{Schw.}}=12a^{2}r^{-6}$, indicating singularity only at the origin, and the same is true also for the trivial extension (2.10). In the nontrivial case (2.11) we have
We can see that for $n=0$ and $n=1$, there is no difference from the Schwarzschild spacetime, $K=K_{\textrm{Schw.}}$.
To analyze the behavior of the Kretschmann scalar, it is convenient to rearrange the relation (3.3) into the form
$\dfrac{1}{12}r^{4}K=\dfrac{1}{\sigma^{2}}\left[1-\dfrac{1}{3}\dfrac{n(n-1)}{(n+1)^{3}}\dfrac{1}{\left(1+\sigma\right)^{2}}\left(\dfrac{3}{4}n+1+(n+1)\sigma\right)\right],$ (3.4)
where $\sigma=r/a$. With the use of the dimensionless radial coordinate $\sigma$, we can depict the behavior of the Kretschmann scalar in two cases with different signs of constant $a$ in one plot. Positive values of $\sigma$ correspond to positive constant $a$, and negative $\sigma$ corresponds to negative $a$ because the radial coordinate $r$ may be only positive. The right-hand side of (3.4) is plotted in Fig. 1 for some numbers of extra dimensions $n$, including the limit $n\to\infty$.
Fig. 1: Right-hand side of (3.4) encoding behaviour of the Kretschmann scalar (3.3) as a function of dimensionless radial coordinate $\sigma$ for various numbers of extra space dimensions.
Unless $a=0$, there is always the central singularity at $r=0$. In the case with positive $a$, there is no additional singularity, but for negative values of constant $a$, the Kretschmann scalar indicates a singularity at the horizon, $r=-a$ or $\sigma=-1$. Other examples of such singularities can be found, for example, in [49, 50].
Another hint of a horizon singularity for negative values of $a$ can be found by investigating radial null geodesics. For an ingoing null geodesic, we can use a parameterization of the radial coordinate as $r=-a+\Delta-\lambda$, where $\Delta$ is a positive constant, and $\lambda$ is the geodesic parameter with values in the interval $\lambda\in[0,\Delta]$. It is sufficient to assume the smallness of $\Delta$ compared to $|a|$. At $\lambda=0$ the ingoing signal starts near the horizon, and at $\lambda=\Delta$ it reaches it. In the case with the trivial extension of the Schwarzschild spacetime (2.10), the amount of the coordinate time needed for a signal to complete this path can be calculated as

### v2

# Properties of generalized Schwarzschild spacetimes

@@ Abstract
We show that an ansatz for $1+3+n$ dimensional static spacetime with spherical symmetry in three dimensions and Euclidean symmetry in $n$ dimensions, parametrized by only one function of radial coordinate, leads to a limited set of vacuum solutions of the Einstein field equations. They can also be identified as Weyl solutions. We investigate properties of these spacetimes through the Kretschmann scalar, Newtonian mass defined through the Newtonian limit, Komar mass, Einstein, Landau–Lifshitz, and ADM mass. In addition to $1+3+n$ dimensional Minkowski spacetime, there are two classes of solutions. The first class is a trivial product of the Schwarzschild spacetime and Euclidean spaces in extra dimensions, while the second class is non-trivial. In the case with no horizon, there is a naked singularity, all masses are equal, and they are negative. In the case when there is a horizon, this horizon accommodates a physical singularity, which corresponds to Kaluza–Klein bubbles featuring exotic properties. Einstein, Landau–Lifshitz, and ADM masses are positive, while Newtonian and Komar masses are negative. This differentiates these solutions from trivial higher-dimensional extensions of the Schwarzschild solution.
@@ 1 Introduction
Schwarzschild spacetime [2] is one of the most prominent examples of vacuum solutions of the Einstein field equations. It can be interpreted as a black hole [3], and such objects are present in our Universe [4, 5, 6]. Black holes generate Hawking radiation [7], providing a laboratory for testing quantum effects in gravity. Modification of the original Schwarzschild solution to an arbitrary number of space dimensions was first found by Tangherlini [8]. Rotating black holes with the Kerr metric [9] can also be generalized to an arbitrary number of space dimensions. They are known as Myers–Perry black holes [10, 11]. There are also other different higher-dimensional generalizations [12], including, for example, black rings [13].
@@
Physical motivation for extra dimensions originates mainly from Kaluza–Klein models [14, 15] attempting to solve the hierarchy problem [16, 17] and string theory [18, 19, 20] developed as a promising candidate for a quantum theory of gravity. These theories usually require some kind of compactification of extra dimensions. Small extra dimensions can be present in a different class of objects known as black strings, p-branes, or black branes [21, 22, 23, 24, 25, 26]. They may have black hole properties in macroscopic dimensions while containing compactified extra dimensions from String theory [27, 28, 29, 30]. Higher-dimensional objects provide valuable testing grounds for studying quantum effects in general relativity [31, 32, 33].
@@
This paper provides an analysis of a special class of vacuum solutions in $1+3+n$ dimensions under the assumption of spherical symmetry in the three-dimensional part and Euclidean symmetry in the $n$-dimensional part, $SO(3)\times E(n)$. Extra dimensions may be compactified into a flat $n$-torus with $S^{1}\times...\times S^{1}$ topology. The number of the standard space dimensions that are not compactified is fixed to three in this work since it is the physically most relevant case. Disregarding $1+3+n$ dimensional Minkowski spacetimes, there will be two different classes of solutions. The first class represents a simple extension of the $1+3$ dimensional Schwarzschild spacetime with the addition of $n$ flat dimensions, with $g_{AB}=\delta_{AB}$, where indices $A$ and $B$ run through $n$ additional dimensions. In contrast, the second class features a non-trivial form of $g_{AB}$ components of the spacetime metric. These solutions can also be found through Weyl solutions [34] and their higher-dimensional generalizations [35] with the non-trivial class corresponding to so-called Kaluza-Klein bubbles [36, 37].
@@
The following section 2 provides vacuum solutions of the Einstein field equations with two classes mentioned above, section 3 includes further analysis of properties of these spacetimes based on the Kretschmann scalar, Newtonian limit, Komar mass, conserved masses from Einstein and Landau–Lifshitz pseudotensors, and ADM mass. Section 4 is dedicated to summary and discussion. The usual convention with the light speed set to unity, $c=1$, together with the spacetime metric signature (,+,+,)(-,+,+,...) will be used throughout this paper. The index 0 will be used for a time coordinate, lowercase Latin indices will indicate three space coordinates, $i=1,2,3$, and capital Latin indices will run through $n$ extra dimensions, A=1,,nA=1,...,n. All spacetime indices will be collectively denoted by Greek letters, $\mu=0,i,A$.
2 Two classes of vacuum solutions
@@
The existence of two different classes of extensions of the Schwarzschild spacetime is obvious in the case with only one extra space dimension. The metric of the trivial case can be written as
@@
where $t$ is time, $a$ is a constant, $r$ is radial coordinate, $d\Omega_{(2)}^{2}=d\vartheta^{2}+\sin^{2}\vartheta d\varphi^{2}$ is line element of a unit $2$-sphere, and $\zeta$ denotes coordinate in the additional dimension. Since the extra dimension is added most trivially, the spacetime with this metric is a vacuum solution like the original Schwarzschild spacetime. The Ricci tensor remains zero if this metric is rewritten through a formal change $\zeta\to it$, $t\to i\zeta$. The resulting metric reads
@@
Metrics (2.1) and (2.2) describe two spacetimes with different properties, which, for example, can be tested by a slowly moving particle far away from the center in the Newtonian limit. Metric (2.1) represents the simplest black string or black ring, and (2.2) is the simplest Kaluza–Klein bubble. This simple example with only one extra space dimension can be generalized to an arbitrary number of extra dimensions, which will be examined in what follows.
Our goal is to find a static vacuum solution of the Einstein field equations with three-dimensional spherical symmetry and $n$-dimensional Euclidean symmetry in extra dimensions, $SO(3)\times E(n)$. This can also be interpreted as a stack of uniform $n$-branes in $1+3+n$ dimensions. We start with the ansatz
@@
where $\alpha$, $\beta$ and $\gamma$ are constants, and $f=f(r)$ is a function of only the radial coordinate. A more general ansatz may be parametrized by two [38] or three independent functions of the radial coordinate [24], but in our analysis, we assume that these functions are given by different powers of one function $f(r)$.
There are two reasons to start with the special choice of ansatz (2.3). A pragmatic reason is that with this choice, the condition of the Ricci tensor being zero reduces to simple algebraic relations. The second reason is that one would expect only one potential event horizon, which occurs only when all components of the spacetime metric $g_{00}$, $g_{rr}$ and $g_{AB}$ are singular at the same radial coordinate, and the simplest way to achieve this is to consider our form of the ansatz. In principle, there may be vacuum solutions with $SO(3)\times E(n)$ symmetry exhibiting more than one horizon; however, with the ansatz (2.3), we will not be able to find them.
@@
$R_{00}=\alpha f^{\alpha-\beta}F_{1},\quad R_{rr}=F_{2},\quad R_{\vartheta\vartheta}=F_{3},$ (2.4)
$R_{\varphi\varphi}=F_{3}\sin^{2}\vartheta,\quad R_{AB}=-\gamma f^{\gamma-\beta}F_{1}\delta_{AB},$
@@
$F_{1}=\dfrac{1}{r}\dfrac{f^{\prime}}{f}+\dfrac{1}{4}\left(\alpha-\beta+n\gamma-2\right)\left(\dfrac{f^{\prime}}{f}\right)^{2}+\dfrac{1}{2}\dfrac{f^{\prime\prime}}{f},$ (2.5)
@@
$F_{3}=1-f^{-\beta}\left[1+\dfrac{1}{2}\left(\alpha-\beta+n\gamma\right)r\dfrac{f^{\prime}}{f}\right],$
with prime denoting differentiation with respect to the radial coordinate. All three functions $F_{1}$, $F_{2}$, and $F_{3}$ must be zero to have the vacuum solution.
@@
where $a$ is an integration constant, and $q$ is given by constants from the ansatz. This solution is invalid if $\alpha-\beta+n\gamma=0$, and this case has to be treated separately. It is easy to see that in this case, the condition $F_{3}=0$ implies $f=1$, and we retrieve the simple $1+3+n$ dimensional Minkowski spacetime. The same is true if the integration constant $a$ is zero. By inserting (2.6) into expressions for $F_{1}$ and $F_{2}$ in (2.5) we find
@@
$\Phi=\dfrac{1}{\left(\alpha-\beta+n\gamma\right)^{2}}\dfrac{ar^{q-2}}{\left(1+ar^{q}\right)^{2}},$ (2.8)
@@
$C_{3}=2\beta^{2}-\alpha^{2}-n\gamma^{2}-\left(\alpha+n\gamma\right)^{2}.$
The problem of finding vacuum solutions of the form (2.3) is now reduced to the system of algebraic equations, $C_{1}=C_{2}=C_{3}=0$.
First two conditions $C_{1}=0$ and $C_{2}=0$ are satisfied if $\alpha=-\beta-n\gamma$, which implies q=1q=-1, and by inserting this expression for $\alpha$ into the relation for $C_{3}$, we also find
@@
The condition $C_{3}=0$ is then satisfied by two sets of possible solutions. One set corresponds to the choice $\gamma=0$, and for the second set, we have $2\beta+(n+1)\gamma=0$. With the choice $\gamma=0$ we have $\alpha=-\beta$. Components of the spacetime metric are then given by $f^{\alpha}=1+ar^{-1}$, $f^{\beta}=\left(1+ar^{-1}\right)^{-1}$, and $f^{\gamma}=1$, and the corresponding spacetime metric can be written as
@@
This is only a trivial extension of the original $1+3$ dimensional Schwarzschild solution, with metric (2.1) as a special case for $n=1$. More interesting solutions correspond to the second choice with $2\beta+(n+1)\gamma=0$, where $\alpha/\beta=(n-1)/(n+1)$, and $\gamma/\beta=-2/(n+1)$. Spacetime metric then can be written as
@@
We can also check that for $n=1$ we retrieve (2.2). In the case with no extra dimensions, $n=0$, both sets of solutions (2.10) and (2.11) automatically reduce to the Schwarzschild solution. From now on, we will address (2.10) as the trivial case and (2.11) as the non-trivial. Solutions (2.10) and (2.11) differ in $g_{00}$ and $g_{AB}$ components. The form of $g_{00}$ determines gravitational redshift as well as the way of fixing the integration constant $a$ by relating it to the gravitating mass in the Newtonian limit $M_{\textrm{New.}}$, while $g_{AB}$ components measure the size of extra dimensions. This will be discussed in more detail in the next section.
Spacetime metric in the non-trivial case (2.11) can be written also in the generalized Weyl form [35],
$ds^{2}=-e^{2A}dt^{2}+e^{2B}d\varphi^{2}+e^{2C}\delta_{AB}d\zeta^{A}d\zeta^{B}+e^{2D}\left(d\rho^{2}+dz^{2}\right),$ (2.12)
where functions $A$, $B$, $C$, and $D$ depend on two space coordinates $\rho$ and $z$, and they are solutions of Laplace’s equation in cylindrical coordinates $(\rho,\varphi,z)$ with infinitely thin rod segments representing sources at $\rho=0$. By choosing these functions as
$A=\frac{1}{2}\frac{1-n}{1+n}\ln\frac{Q+a}{Q-a},\quad B=\ln\rho-\frac{1}{2}\ln\frac{Q+a}{Q-a},$ (2.13)
$C=\frac{1}{1+n}\ln\frac{Q+a}{Q-a},\quad D=\frac{1}{2}\ln\frac{(L-a)^{2}}{4Q_{+}Q_{-}},$
where $Q=Q_{+}+Q_{-}$, $Q_{\pm}=\sqrt{(z\pm a/2)^{2}+\rho^{2}}$, and by performing the coordinate transformation $z=r(1+a/(2r))\cos\vartheta$, $\rho=r\sqrt{1+a/r}\sin\vartheta$, we can rewrite (2.12) into the form (2.11).
Different solutions similar to (2.11) can be found in the literature. Authors of [21] worked with the low-energy string theory action in ten dimensions, and they found a brane solution with a single Ramond–Ramond charge with non-trivial $g_{AB}$ components in terms of our notations. This has been studied further in many other works [39, 40, 41, 42, 43, 44]. Extra dimensions parametrized by coordinates $\zeta^{A}$ in (2.11) may be macroscopic with $\mathbb{R}^{n}$ topology, microscopic with $n$-torus topology, or a combination of both cases with some macroscopic and some microscopic dimensions. In all cases, there is Euclidean symmetry in extra dimensions, but one can assign extra dimensions to the part that is only spherically symmetric [45]. Note also that considering extra time dimensions instead of extra space dimensions, or having mixed time and space extra dimensions, would not change the results because changing the signs of $g_{AB}$ components of the metric keeps the Ricci tensor zero. Compactification of the time dimension may also be considered, with implications for branes [46].
We will study some properties of the found solutions (2.10) and (2.11) in the following section. To focus on physically most relevant cases, let us assume that all extra dimensions are microscopic with $S^{1}\times...\times S^{1}$ topology of a flat $n$-torus. Without loss of generality, we may consider all extra coordinates $\zeta^{A}$ to be in the interval $\zeta^{A}\in[0,L)$, with the same periodicity $L$ for every $\zeta^{A}$.
@@ 3 Physical properties
In the case of the ordinary Schwarzschild metric, there is no physical singularity on the horizon. This can be easily revealed by the calculation of the Kretschmann scalar invariant defined through the Riemann curvature tensor as $K=R^{\mu\nu\lambda\sigma}R_{\mu\nu\lambda\sigma}$. For the Schwarzschild spacetime, it is $K_{\textrm{Schw.}}=12a^{2}r^{-6}$, indicating a singularity only at the origin, and the same is true also for the trivial extension (2.10). In the non-trivial case (2.11) we have
We can see that for $n=0$ and $n=1$, there is no difference from the Schwarzschild spacetime, $K=K_{\textrm{Schw.}}$. Unless $a=0$, there is always the central singularity at $r=0$. In the case with positive $a$, there is no additional singularity, but for negative values of constant $a$, the Kretschmann scalar indicates a singularity at the horizon, r=ar=-a, for $n\geq 2$. Other examples of such singularities can be found, for example, in [47, 48].
Since there is a horizon which is not associated with any physical singularity in the Schwarzschild solution, it is possible to find its maximal analytic extension [49]. Obviously, the same is true also for the trivial case with extra dimensions (2.10) if $a$ is negative. In other cases, there is no possibility to construct the maximal extension either because of the lack of horizon if $a$ is positive, or because of the presence of the physical singularity at the horizon. This also means that two regions inside and outside the spherical singularity represent two separate spacetimes. The spacetime corresponding to the outer region has properties of a Kaluza–Klein bubble [36, 37], since from the point of view of the outer region, there is no spacetime below the spherical singularity.
The Newtonian limit provides a very important insight into properties of branes containing spherically symmetric $1+3$ dimensional part [50, 51]. We will examine solutions (2.10) and (2.11) in this limit under the assumption of $L$-periodicity of coordinates $\zeta^{A}$ in extra dimensions. Macroscopic objects of size larger than $L$ can move only in the $1+3$ dimensional section. The same is true for sufficiently long-wavelength photons or particles with large enough de Broglie wavelengths. From the perspective of such objects, the spacetime in regions far from the origin, $r\gg|a|$, can be approximated as
@@
where $\phi$ is the Newtonian potential. Without any extra microscopic dimensions, it can be shown that $\psi=\phi$ in harmonic coordinates. The geodesic equation for a slowly moving object reduces to the Newtonian equation of motion, $\ddot{x}^{i}\approx-\phi_{,i}$, in this limit.
In order to examine the Newtonian limit for spacetimes with metrics (2.10) and (2.11), we have to replace the radial coordinate $r$ with the coordinate $\rho$ defined by the transformation $r=\rho\left(1-a/(4\rho)\right)^{2}$. This new coordinate is usually called the isotropic radial coordinate because in the case of $1+3$ dimensional Schwarzschild metric, the space part of the metric becomes conformally Euclidean. In our case, this is true only for the three-dimensional and $n$-dimensional sections separately. For positive values of $a$, the isotropic radial coordinate $\rho$ covers the entire spacetime except for the central singularity twice, while for negative $a$, it covers twice only the region above the horizon. To avoid the double coverage, we may consider $\rho$ only in the interval $|a|/4,\infty)$. In the trivial case (2.10) we have
where functions $\chi_{+}$ and $\chi_{-}$ are defined as $\chi_{\pm}=1\pm a/(4\rho)$, and space coordinates $x^{i}$ are Cartesian such that $(x^{1})^{2}+(x^{2})^{2}+(x^{3})^{2}=\rho^{2}$. This leads to
$ds^{2}\approx-\left(1+\dfrac{a}{\rho}\right)dt^{2}+\left(1-\dfrac{a}{\rho}\right)\delta_{ij}dx^{i}dx^{j}+\delta_{AB}d\zeta^{A}d\zeta^{B},$ (3.4)
in the Newtonian limit, $\rho\gg|a|$, with the form of the Newtonian potential, $\phi=(a/2)\rho^{-1}$, and the usual relation between the constant $a$ and the mass of the gravitating body in the Newtonian theory, $a=-2\kappa M_{\textrm{New.}}$, where $\kappa$ denotes the Newtonian gravitation constant. In the non-trivial case (2.11), the spacetime metric can be rewritten as
$ds^{2}=-\left(\dfrac{\chi_{+}}{\chi_{-}}\right)^{-2\frac{n-1}{n+1}}dt^{2}+\left(\chi_{-}\right)^{4}\delta_{ij}dx^{i}dx^{j}+\left(\dfrac{\chi_{+}}{\chi_{-}}\right)^{\frac{4}{n+1}}\delta_{AB}d\zeta^{A}d\zeta^{B}\approx$ (3.5)